OpenStudy (lgbasallote):

obtain the particular solution satisfying the initial condition indicated \[y' = x \text{exp} (y - x^2)\] when x = 0; y = 0 steps please?

OpenStudy (experimentx):

let's try using variable separation ... \[ e^{y - x^2} = \frac{e^y}{e^{x^2}}\]

OpenStudy (lgbasallote):

where did e^y and e^x^2 come from?

OpenStudy (experimentx):

\[ \huge \frac{dy}{dx} = \frac{x e^y}{e^{x^2}}\] \[ \huge \int \frac{dy}{e^y} = \int \frac{x}{e^{x^2}} dx\]

OpenStudy (experimentx):

x^(a+b) = x^a * x^b

OpenStudy (lgbasallote):

i see..i think that will be all thanks

OpenStudy (experimentx):

yw